OpenStudy (1018):

Please help. 4 derivative problems.

OpenStudy (1018):

Here's the first one: \[\frac{ \sqrt[3]{x} }{ 1+x }\]

OpenStudy (unklerhaukus):

use the quotient rule \[\left(\frac{u}{v}\right)'=\frac{u'v-uv'}{v^2}\]

OpenStudy (unklerhaukus):

Identify \(u\) and \(v\), then find \(u'\), \(v'\), and \(v^2\)

OpenStudy (1018):

Here's what I got so far. Could you kindly check. \[\frac{ (1+x)\frac{ 1 }{ 3 } (x)^{\frac{ -2 }{ 3 }} - (\sqrt[3]{x}) }{ (1+x)^{2} }\]

OpenStudy (unklerhaukus):

yes, that is correct, now simplify

OpenStudy (1018):

Ok, thanks! Please help me simplify though. Haha. I'm not sure if I'd be right.

OpenStudy (unklerhaukus):

simplify this bit first \[(1+x)\frac{ 1 }{ 3 } (x)^{\frac{ -2 }{ 3 }}\]

OpenStudy (1018):

\[\frac{ x ^{\frac{ -2 }{ 3 }} + x ^{\frac{ 1 }{ 3 }} }{ 3 }\]

OpenStudy (1018):

How's that?

OpenStudy (unklerhaukus):

good, now combine that other term in the numerator

OpenStudy (1018):

Oh, yeah. So it's x^-1/3 ?

OpenStudy (unklerhaukus):

um \(-\sqrt[3]x\) is equal to \(-x^{1/3}\)

OpenStudy (unklerhaukus):

\[\frac{ (1+x)\frac{ 1 }{ 3 } (x)^{\frac{ -2 }{ 3 }} - (\sqrt[3]{x}) }{ (1+x)^{2} }\] \[\frac{ \frac{ x ^{-2/3} + x ^{1/3} }{ 3 } -x^{1/3} }{ (1+x)^{2} }\]

OpenStudy (1018):

Oh wait, back up. Haha. My last answer was I just combined the numerator.

OpenStudy (1018):

With the Xs, with the exponent, the one with the 3 denominator.

mathslover (mathslover):

Uncle, what about substitution ?

OpenStudy (1018):

Guys, I'll close this now. I'll just stick to my answer. Hope it's right. lol. Thanks!

mathslover (mathslover):

http://www.wolframalpha.com/input/?i=differentiate+cube+root+of+x+%2F+%281%2Bx%29+ This is what wolfram gives as the answer.

OpenStudy (unklerhaukus):

ah so we are close to the answer, we just to finish the simplification (rearrangement )

OpenStudy (unklerhaukus):

\[\frac{ \frac{ x ^{-2/3} + x ^{1/3} }{ 3 } -x^{1/3} }{ (1+x)^{2} }\] \[\frac{x^{-2/3} + x ^{1/3} -3x^{1/3} }{ 3(1+x)^{2} }\]\[...\]

mathslover (mathslover):

If we use Substitution : cube root of x = u diff. both sides ... \(dx = \cfrac{du}{\cfrac{1}{3} x^{-2/3} } \) \(\cfrac{d}{dx} \left( \cfrac{u}{1+u^3} \right) \times \cfrac{1}{3}x^{-2/3} \) \(\cfrac{(1+x) - x^{1/3} (3 \times x^{2/3} ) }{(x+1)^2} \times \cfrac{1}{3} x^{-2/3} \) \(\cfrac{1-2x}{3(x+1)^2 \times x^{2/3} } \)

OpenStudy (unklerhaukus):

\[\frac{x^{-2/3} + x ^{1/3} -3x^{1/3} }{ 3(1+x)^{2} }\] \[\frac{x^{-2/3} -2x ^{1/3}}{ 3(1+x)^{2} }\] \[\frac{(1 -2x)x^{-2/3}}{ 3(1+x)^{2} }\] \[\frac{(1 -2x)}{ 3(1+x)^{2}x^{2/3} }\] the answers all match yayayay\(\checkmark\checkmark\checkmark\)

mathslover (mathslover):

Well Done Mr. Rhaukus :P